The number of common prime factors of 75, 60, 105 is (A) 2 (B) 3 (C) 4 (D) 5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student
by prime factorization
75 = 3 × 5 × 5
60 = 2 × 2 × 3 × 5
105 = 3 × 5 × 7
Thus, common prime factors in the given three numbers are 3 and 5.
Hence, the number of common prime factors of 75, 60, 105 is 2.

Thanks